A combined total of $20,000 is invested in two bonds that pay 8% and 9.5% simple intrest. The annual interest is $1675. How much is invested in each bond?
Answer by checkley71(8403)

.095X+.08(20000-X)=1675
.095X+1600-.08X=1675
.015X=1675-1600
.015X=75
X=75/.015
X=5000 INVESTED @ 9.5%
20000-5000=15000 INVESTED @ 8%
PROOF
.095*5000=.08*15000=1675
475+1200=1675
1675=1675